# A quantum random access memory (QRAM) using a polynomial encoding of binary strings

### v1

@@ Abstract
Quantum algorithms claim significant speedup over their classical counterparts for solving many problems. An important aspect of many of these algorithms is the existence of a quantum oracle, which needs to be implemented efficiently in order to realize the claimed advantages in practice. A quantum random access memory (QRAM) is a promising architecture for realizing these oracles. In this paper we develop a new design for QRAM and implement it with Clifford+T circuit. We focus on optimizing the T-count and T-depth since non-Clifford gates are the most expensive to implement fault-tolerantly in most error correction schemes.
Integral to our design is a polynomial encoding of bit strings and so we refer to this design as $\text{QRAM}_{poly}$. Compared to the previous state-of-the-art bucket brigade architecture for QRAM, we achieve an exponential improvement in T-depth, while reducing T-count and keeping the qubit-count same. Specifically, if $N$ is the number of memory locations to be queried, then $\text{QRAM}_{poly}$ has T-depth $O(\log\log N)$, T-count $O(N-\log N)$ and uses $O(N)$ logical qubits, while the bucket brigade circuit has T-depth $O(\log N)$, T-count $O(N)$ and uses $O(N)$ qubits.
Combining two $\text{QRAM}_{poly}$ we design a quantum look-up-table, $\text{qLUT}_{poly}$, that has T-depth $O(\log\log N)$, T-count $O(\sqrt{N})$ and qubit count $O(\sqrt{N})$. A quantum look-up table (qLUT) or quantum read-only memory (QROM) has restricted functionality than a QRAM. For example, it cannot write into a memory location and the circuit needs to be compiled each time the contents of the memory change. The previous state-of-the-art CSWAP architecture has T-depth $O(\sqrt{N})$, T-count $O(\sqrt{N})$ and qubit count $O(\sqrt{N})$. Thus we achieve a double exponential improvement in T-depth while keeping the T-count and qubit-count asymptotically same. Additionally, with our polynomial encoding of bit strings, we develop a method to optimize the Toffoli-count of circuits, specially those consisting of multi-controlled-NOT gates.
@@ 1 Introduction
Quantum computers hold immense promise to outperform classical computers in many applications. Over the years numerous quantum algorithms have been developed that claim speedups over their classical counterparts in various problems, for example, unstructured database search [Gro96], optimization [vAGGdW20], quantum chemistry algorithms [RWS+17, BWM+18, BBMC20, CRO+19, RBM+23], data processing for machine learning [dO09, BWP+17, HHL09, HKT+22, CHI+18, AAD+15, BDLK19], cryptography [Sho94, Kup13], image processing [OCKK21]. Many of these algorithms require access to oracles in order to fetch classical data and in practice, this is a non-trivial task. It is important to specify the details of implementations of these oracles in order to claim a genuine quantum speedup [Aar15]. Efficient implementation of oracles can reduce the crossover of runtime between classical and quantum advantage from years to days [vBLH+21].
Till date, the most general-purpose design for the implementation of quantum oracles is a quantum random access memory (QRAM) [GLM08a, KMR05, Ble10, LWS+23, PCG23], which analogous to a classical random access memory (RAM), returns the element stored in a particular memory location. Specifically, suppose there are $N$ memory locations, each indexed by an integer $i\in\{0,1,\ldots,N-1\}$ and element $x_{i}$ is stored in location $i$. Then on input $i$, a classical RAM returns $x_{i}$. This procedure is called “reading” from the memory. A classical RAM is also able to “write” a particular data $x_{i}$ into memory location $i$. With a QRAM we are able to query a superposition of addresses. Let $A$ be the input qubit register storing the memory address to be queried and $B$ be the output register. If |$\ket{\psi_{in}}$ and |$\ket{\psi_{out}}$ are the input and output states, then
@@
and after the operation $x_{i}$ is XOR-ed into the memory location $i$. The oracles described in many algorithms [KP17] do require the writing operation. Giovannetti, Lloyd and Maccone introduced the fanout and bucket-brigade architectures for QRAM in their pioneering work in [GLM08a, GLM08b]. Since then much work has been done to study these designs and improve upon them. Out of the two desgins the bucket-brigade QRAM has become the most popular because it has better noise resilience [AGJO+15, HLGJ21] and fault-tolerant resource estimates [DMGM20]. Several proposals for the experimental implementations of QRAM have been put forth [GLM08a, HXZ+12, MM16, HZZ+19, CDEH+21, Pla22, WPG24], each utilizing the bucket brigade architecture. In [XHF+23] the authors propose a design implementing an $n$-bit QRAM on hardware nominally supported only on an $m$-bit query, where $m<n$. Over the years there have been proposals for various implementations of QRAM using different techniques, often for specific applications [ADMG+16, SRWDM18, JS19, PPR19, DVDAPDS20, ASY21, OCKK21, ZAAK+21, ASY23, BCC+23, CAY23, CXW+23, PLG23, LHJ23, LJ24, DH24, Hun24], and thus QRAMs have been used for a wide variety of tasks like neural networks, data processing, quantum communication, image processing, cryptanalysis, quantum simulation, circuit synthesis, state preparation, etc.
@@
A circuit implementing a QRAM needs to be compiled and optimized only once, while the contents of the memory are free to change. But it has the disadvantage of a significant space overhead. A bucket-brigade QRAM for $N$ memory locations require $O(N)$ T gates, $O(N)$ ancillae and has T-depth $O(\log N)$ [DMGM20]. In order to reduce the number of ancillae many algorithms use a sequence of multi-controlled-NOT gates, also known as quantum read-only memory (QROM) [BWM+18, DMGM20, RBM+23, MSW24]. This can be implemented with $O(N)$ T gates, $O(\log N)$ ancillae and $O(N)$ T-depth. Inspite of a lower qubit count, one disadvantage of a QROM is the exponentially higher T-depth which is not desirable for an efficient fault-tolerant implementation. Another disadvantage of QROM is the fact that we need to know the contents of the memory in advance. Each time the database changes, the circuit needs to be recompiled and optimized. There are other architectures, as in [PPR19, DVDAPDS20], that perform queries in $O(N\log N)$ time using $O(\log N)$ qubits.
@@
Many hybrid architectures have been proposed that interpolate between these two extremes and leverage their space-time tradeoff [BGM+19, DMGM20, HKRS22, KSRZ22, LKS24, ZSL24, MSW24]. Notable among these is the CSWAP architecture [LKS24], which can be thought of as a combination of a QROM and a specific QRAM. It has T-count $O(\sqrt{N})$, number of ancillae $O(\sqrt{N})$ and T-depth $O(\sqrt{N})$. Here we mention that in literature the QRAM, QROM and these hybrid architectures are also used to build quantum look-up-table (qLUT) and so the names are often used interchangeably. These are required to perform restricted tasks. The contents of the table or database are known and this can be leveraged to design circuits with better resource estimates like T-count.
@@ 1.1 Our contributions
In this paper we propose an architecture for a QRAM, mainly aimed at reducing the non-Clifford gate complexity of the circuit implementation. We implement our circuits with the fault-tolerant, universal Clifford+T gate set because it implements more unitaries exactly compared to other fault-tolerant, universal gate sets [MM21, Muk24a, Muk24c, Muk24b]. In most error correction schemes the cost of implementing the non-Clifford T gate dwarfs the cost of all other gates by factors of 100 or greater [FMMC12]. Thus it is important to optimize the number of T-gates or T-count. It is also important to optimize the T-depth [Fow12, FMMC12, ADMG+16, HS20, DMGM20, GMM22a], which is related to the running time. A T-depth-1 corresponds to a set of T gates that can be implemented in parallel. We also refer to it as a layer or stage. So a T-depth $\mathcal{T}_{d}$ for a circuit implies $\mathcal{T}_{d}$ such stages, where in each stage the T gates are implemented in parallel. The product of T-depth and number of logical qubits is taken as a parameter to measure the rough cost of fault-tolerant implementation in the surface code [ADMG+16, DMGM20, LKS24]. Our contributions in this paper can be summarized in the following points.
@@
Bucket-brigade [GLM08a, DMGM20] $O(\log N)$ $O(N)$ $O(N)$
CSWAP [LKS24] $O(\sqrt{N})$ $O(\sqrt{N})$ $O(\sqrt{N})$
@@
(I) Suppose we aim to construct an $n$-qubit QRAM, that is, the number of memory locations is $N=2^{n}$. We describe an encoding by which any $n$-bit string $(b_{1},b_{2},\ldots,b_{n})$ can be represented by a polynomial over $n$ Boolean variables $\{x_{1},x_{2},\ldots,x_{n}\}$. The set of $N$ polynomials are such that for any given bit string (or values for these variables), only one of these polynomials has value 1 mod 2 and the rest have 0 mod 2. We call these the encoding polynomials and they have a 1-to-1 correspondence with the bit strings. Now, the $N$ memory locations are each indexed by an $n$-bit string, which we refer to as the address. So there is an encoding polynomial associated with each memory location (Section 2).
(II) With the help of the encoding polynomials we develop a quantum random access memory, which we call $\text{QRAM}_{poly}$ (Section 3). We allocate $N$ ancillae qubits, each selecting a particular memory location. These are initialized to the |$\ket{0}$ state and we index them with the addresses of the locations they select. In each ancilla, using Toffolis and CNOTs, we implement the encoding polynomial corresponding to its index. Thus, for a particular input address, corresponding to a certain value of the variables or a certain bit string, one of these ancillae flips to |$\ket{1}$. Then we can perform any operation on the corresponding memory location, controlled on this ancilla.
We show that we require $N-\log N-1$ Toffoli gates to implement a circuit. We can parallelize this circuit, using an additional $O(N)$ ancillae in order to achieve a Toffoli-depth of $\log\log N$. This implies a T-count of $O(N-\log N-1)$ and T-depth of $O(\log\log N)$. Thus compared to previous bucket-brigade architecture [DMGM20] we achieve an exponential improvement in the T-depth, reduce the T-count, while keeping the number of ancillae nearly the same.
(III) We use two $\text{QRAM}_{poly}$ to design a QROM or qLUT. This is a hybrid architecture and we achieve a T-count of $O(\sqrt{N})$, T-depth $O(\log\log N)$ and ancillae count of $O(\sqrt{N})$. Thus, here we achieve a double exponential improvement in T-depth compared to previous designs [LKS24, ZSL24], while keeping the T-count and ancillae count asymptotically similar. We refer to this desgin as $\text{qLUT}_{poly}$ (Section 4.1). In Table 1 we have summarized and compared the cost of implementation of our $\text{QRAM}_{poly}$ and $\text{qLUT}_{poly}$ with some previous works.
(IV) The encoding polynomials can also be used to optimize the Toffoli-count of circuits, dominated by multi-controlled-NOT gates. We describe a procedure TOFFOLI-OPT-POLY for performing such compiling in Section 4.2.
@@ 3 Quantum random access memory ($\text{QRAM}_{poly}$)
We perform the following steps.
@@
    Select and compute : Using a Toffoli controlled on $a_{j}$ and $M_{j}$ we copy the memory content onto the output bus (Figure 1(b)). This is equivalent to “reading” from the memory. For “writing” into the memory we reverse the control and target at the output bus and $M_{j}$. We can also do other operations, for example, applying phase on selected memory locations, as required in Grover’s algorithm. In this case we do not require the Toffolis, as shown in Figure 1(b). Instead, we use CZ on each memory location where the control is on the selecting ancilla.
@@ 3.1 Cost of implementation
Now we discuss the cost of implementing a circuit with the Clifford+T gate set. We focus on optimizing the non-Clifford resource, that is, the T-count and T-depth, as discussed earlier in Section 1.1. We first bound the Toffoli depth and number of compute-uncompute Toffoli pairs. In literature there are different implementations of a Toffoli gate with the Clifford+T gate set, some optimizing its T-count, while there are others that optimize T-depth. The circuit in [Jon13, Gid18] gives the lowest T-count of 4 and has a T-depth of 2. It uses logical AND gadget which does the multiplication. One advantage of this circuit is the fact that if we have a compute-uncompute pair then the uncomputation part does not require any T-gate, but it uses classical measurements. Another circuit is the one in [Sel13] which has a T-count of 7, T-depth of 1 and uses 4 extra ancillae. Depending on whichever parameter we want to optimize, one implementation can be favoured over the other.
We have separated the cost after the computation of the encoding polynomials since the cost of this part can change depending upon the required operations, as discussed in Step 3.

### v2

@@ Abstract
Quantum algorithms claim significant speedup over their classical counterparts for solving many problems. An important aspect of many of these algorithms is the existence of a quantum oracle, which needs to be implemented efficiently in order to realize the claimed advantages in practice. A quantum random access memory (QRAM) is a promising architecture for realizing these oracles. In this paper we develop a new design for QRAM and implement it with Clifford+T circuit. We focus on optimizing the T-count and T-depth since non-Clifford gates are the most expensive to implement fault-tolerantly in most error correction schemes. Integral to our design is a polynomial encoding of bit strings and so we refer to this design as $\text{QRAM}_{poly}$. Compared to the previous state-of-the-art bucket brigade architecture for QRAM, we achieve an exponential improvement in T-depth, while reducing T-count and keeping the qubit-count same. Specifically, if $N$ is the number of memory locations to be queried, then $\text{QRAM}_{poly}$ has T-depth $O(\log\log N)$, T-count $O(N-\log N)$ and uses $O(N)$ logical qubits, while the bucket brigade circuit has T-depth $O(\log N)$, T-count $O(N)$ and uses $O(N)$ qubits. Combining two $\text{QRAM}_{poly}$ we design a quantum look-up-table, $\text{qLUT}_{poly}$, that has T-depth $O(\log\log N)$, T-count $O(\sqrt{N})$ and qubit count $O(\sqrt{N})$. A quantum look-up table (qLUT) or quantum read-only memory (QROM) has restricted functionality than a QRAM. For example, it cannot write into a memory location and the circuit needs to be compiled each time the contents of the memory change. The previous state-of-the-art CSWAP architecture has T-depth $O(\sqrt{N})$, T-count $O(\sqrt{N})$ and qubit count $O(\sqrt{N})$. Thus we achieve a double exponential improvement in T-depth while keeping the T-count and qubit-count asymptotically same. Additionally, with our polynomial encoding of bit strings, we develop a method to optimize the Toffoli-count of circuits, specially those consisting of multi-controlled-NOT gates.
@@ 1 Introduction
Quantum computers hold immense promise to outperform classical computers in many applications. Over the years numerous quantum algorithms have been developed that claim speedups over their classical counterparts in various problems, for example, unstructured database search [1], optimization [2], quantum chemistry algorithms [3, 4, 5, 6, 7], data processing for machine learning [8, 9, 10, 11, 12, 13, 14], cryptography [15, 16], image processing [17]. Many of these algorithms require access to oracles in order to fetch classical data and in practice, this is a non-trivial task. It is important to specify the details of implementations of these oracles in order to claim a genuine quantum speedup [18]. Efficient implementation of oracles can reduce the crossover of runtime between classical and quantum advantage from years to days [19].
Till date, the most general-purpose design for the implementation of quantum oracles is a quantum random access memory (QRAM) [20, 21, 22, 23, 24], which analogous to a classical random access memory (RAM), returns the element stored in a particular memory location. Specifically, suppose there are $N$ memory locations, each indexed by an integer $i\in\{0,1,\ldots,N-1\}$ and element $x_{i}$ is stored in location $i$. Then on input $i$, a classical RAM returns $x_{i}$. This procedure is called “reading” from the memory. A classical RAM is also able to “write” a particular data $x_{i}$ into memory location $i$. With a QRAM we are able to query a superposition of addresses. Let $A$ be the input qubit register storing the memory address to be queried and $B$ be the output register. If |$\ket{\psi_{in}}$ and |$\ket{\psi_{out}}$ are the input and output states, then
@@
and after the operation $x_{i}$ is XOR-ed into the memory location $i$. The oracles described in many algorithms [25] do require the writing operation. Giovannetti, Lloyd and Maccone introduced the fanout and bucket-brigade architectures for QRAM in their pioneering work in [20, 26]. Since then much work has been done to study these designs and improve upon them. Out of the two desgins the bucket-brigade QRAM has become the most popular because it has better noise resilience [27, 28] and fault-tolerant resource estimates [29]. Several proposals for the experimental implementations of QRAM have been put forth [20, 30, 31, 32, 33, 34, 35], each utilizing the bucket brigade architecture. In [36] the authors propose a design implementing an $n$-bit QRAM on hardware nominally supported only on an $m$-bit query, where $m<n$. Over the years there have been proposals for various implementations of QRAM using different techniques, often for specific applications [37, 38, 39, 40, 41, 42, 17, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], and thus QRAMs have been used for a wide variety of tasks like neural networks, data processing, quantum communication, image processing, cryptanalysis, quantum simulation, circuit synthesis, state preparation, etc.
@@
A circuit implementing a QRAM needs to be compiled and optimized only once, while the contents of the memory are free to change. But it has the disadvantage of a significant space overhead. A bucket-brigade QRAM for $N$ memory locations require $O(N)$ T gates, $O(N)$ ancillae and has T-depth $O(\log N)$ [29]. In order to reduce the number of ancillae many algorithms use a sequence of multi-controlled-NOT gates, also known as quantum read-only memory (QROM) [4, 29, 7, 54]. This can be implemented with $O(N)$ T gates, $O(\log N)$ ancillae and $O(N)$ T-depth. Inspite of a lower qubit count, one disadvantage of a QROM is the exponentially higher T-depth which is not desirable for an efficient fault-tolerant implementation. Another disadvantage of QROM is the fact that we need to know the contents of the memory in advance. Each time the database changes, the circuit needs to be recompiled and optimized. There are other architectures, as in [40, 41], that perform queries in $O(N\log N)$ time using $O(\log N)$ qubits.
@@
Many hybrid architectures have been proposed that interpolate between these two extremes and leverage their space-time tradeoff [55, 29, 56, 57, 58, 59, 54]. Notable among these is the CSWAP architecture [58], which can be thought of as a combination of a QROM and a specific QRAM. It has T-count $O(\sqrt{N})$, number of ancillae $O(\sqrt{N})$ and T-depth $O(\sqrt{N})$. Here we mention that in literature the QRAM, QROM and these hybrid architectures are also used to build quantum look-up-table (qLUT) and so the names are often used interchangeably. These are required to perform restricted tasks. The contents of the table or database are known and this can be leveraged to design circuits with better resource estimates like T-count.
@@ 1.1 Our contributions
In this paper we propose an architecture for a QRAM, mainly aimed at reducing the non-Clifford gate complexity of the circuit implementation. We implement our circuits with the fault-tolerant, universal Clifford+T gate set because it implements more unitaries exactly compared to other fault-tolerant, universal gate sets [60, 61, 62, 63]. In most error correction schemes the cost of implementing the non-Clifford T gate is significantly higher than the Clifford gates. Thus it is important to optimize the number of T-gates or T-count. It is also important to optimize the T-depth [65, 64, 37, 66, 29, 67], which is related to the running time. A T-depth-1 corresponds to a set of T gates that can be implemented in parallel. We also refer to it as a layer or stage. So a T-depth $\mathcal{T}_{d}$ for a circuit implies $\mathcal{T}_{d}$ such stages, where in each stage the T gates are implemented in parallel. The product of T-depth and number of logical qubits is taken as a parameter to measure the rough cost of fault-tolerant implementation in the surface code [37, 29, 58]. Our contributions in this paper can be summarized in the following points.
@@
Bucket-brigade [20, 29] $O(\log N)$ $O(N)$ $O(N)$
CSWAP [58] $O(\sqrt{N})$ $O(\sqrt{N})$ $O(\sqrt{N})$
@@
(I) We develop a quantum random access memory, which we call $\text{QRAM}_{poly}$ (Section 3), with the help of a polynomial encoding of bit strings (Section 2). We show that $\text{QRAM}_{poly}$ can be implemented with $N-\log N-1$ Toffoli gates. We can parallelize this circuit, using an additional $O(N)$ ancillae in order to achieve a Toffoli-depth of $\log\log N$. This implies a T-count of $O(N-\log N-1)$ and T-depth of $O(\log\log N)$. Thus compared to previous bucket-brigade architecture [29] we achieve an exponential improvement in the T-depth, reduce the T-count, while keeping the number of ancillae nearly the same.
(II) We use two $\text{QRAM}_{poly}$ to design a QROM or qLUT. This is a hybrid architecture and we achieve a T-count of $O(\sqrt{N})$, T-depth $O(\log\log N)$ and ancillae count of $O(\sqrt{N})$. Thus, here we achieve a double exponential improvement in T-depth compared to previous designs [58, 59], while keeping the T-count and ancillae count asymptotically similar. We refer to this desgin as $\text{qLUT}_{poly}$ (Section 4.1). In Table 1 we have summarized and compared the cost of implementation of our $\text{QRAM}_{poly}$ and $\text{qLUT}_{poly}$ with some previous works.
(III) The encoding polynomials that are integral to our constructions of QRAM and qLUT have other applications. For example, in Section 4.2 we describe a procedure (TOFFOLI-OPT-POLY) to optimize the Toffoli-count of circuits. Later we also discuss some other potential applications and hence these polynomials may be of independent interest.
@@ 3 Quantum random access memory ($\text{QRAM}_{poly}$)
We perform the following steps. In Figure 1 we have shown a flowchart depicting these steps.
@@
    Select and compute : Using a Toffoli controlled on $a_{j}$ and $M_{j}$ we copy the memory content onto the output bus (Figure 2(b)). This is equivalent to “reading” from the memory. For “writing” into the memory we reverse the control and target at the output bus and $M_{j}$.
@@
Figure 1: (a) A flowchart showing the procedures in $\text{QRAM}_{poly}$. In the figure “I/p” and “enc. poly.” represents “Input” and “encoding polynomials”, respectively. (b) The procedure of evaluating monomials using Toffoli.
@@
3.1 Illustration : Application of $\text{QRAM}_{poly}$ in Grover’s algorithm
To further illustrate the application of $\text{QRAM}_{poly}$ we consider the Grover’s algorithm [1], which gives a thoeretical quadratic speedup over classical search algorithms in an unstructured database. Suppose there are $N$ items in the database. The Grover’s algorithm consists of the following steps. First we initialize the system in the state |$\ket{\Psi}=\ket{0}^{n}$. Then we perform a number of iterations of the following procedure.
  1. 1.
    Set all qubits into an equal superposition state |$\ket{s}$.
    $\text{H}^{\otimes n}\ket{0}^{n}$ $=$ $\frac{1}{\sqrt{N}}\sum_{i=0}^{N-1}\ket{i}=\ket{s}.$
  2. 2.
    Phase-tag the states that represent the values to be searched.
  3. 3.
    Implement a diffusion operator $U_{d}=2\ket{s}\bra{s}-\mathbb{I}$ that amplifies the amplitudes for measuring the states that need to be searched.
At the end of all iterations of the algorithm we perform a measurement in the computational basis. The searched items can be be found by identifying the distinct peaks in the distribution of the measured results.
Steps 2 (phase-tagging) and 3 (diffusion operator) correspond to two successively performed reflections, and thus together they perform a rotation in a 2D-plane. Thus in each iteration of the Grover’s algorithm the state |$\ket{s}$ is rotated closer to a state |$\ket{k}$, that represents a value to be searched. After an optimal number of iterations |$\ket{s}$ is rotated the closest to |$\ket{k}$. Searching one item in an unstructured database with $N$ items requires at most $O\left(\sqrt{N}\right)$ iterations. Classically this search can be done in $O(N)$ time complexity. Thus it is possible to achieve a quadatic speed-up, provided each iteration is done efficiently, or simply put, time complexity of each iteration is considerably less than the number of iterations.
Usually, in theoretical analyses of Grover’s algorithm we assume the existence of a phase-tagging oracle that performs step 2. This oracle has the following functionality :
$O\ket{i}=(-1)^{f(i)}\ket{i}\quad\text{with}\quad f(i)=1\quad\text{if}\quad i\in\{k\}\quad\text{else}\quad f(i)=0.$
An efficient implementation of this oracle is essential in order to achieve the claimed speedup of the Grover’s algorithm in practice. Our $\text{QRAM}_{poly}$ can be used to implement this oracle. Specifically, after we compute the encoding polynomials in order to select a memory location (Step 3) we do not require the Toffolis, as shown in Figure 2(b). Instead, we use CZ on each memory location where the control is on the selecting ancilla. In this way, we apply phase on selected memory locations.
We are not going into more detail of an optimal fault-tolerant implementation of Grover’s algorithm in order to achieve a practical quantum speed-up as this is a stand-alone research topic [73] and beyond the scope of this paper. But briefly we want to summarize this section by emphasizing that an efficient implementation of the phase-tagging oracle is crucial for the practical speed-up of Grover’s algorithm. A QRAM (with proper modification) can be used and faster this QRAM, the better it is. In later sections we give a detail analysis of the cost of fault-tolerant implementation of $\text{QRAM}_{poly}$ (though with read/write operations) with the surface code and show that it is much faster and consumes less number of qubits, compared to other QRAMs. Thus it can also be used to have a faster fault-tolerant implementation of Grover’s algorithm.
@@ 3.2 Cost of implementation
Now we discuss the cost of implementing a circuit with the Clifford+T gate set. We focus on optimizing the non-Clifford resource, that is, the T-count and T-depth, as discussed earlier in Section 1.1. We first bound the Toffoli depth and number of compute-uncompute Toffoli pairs. In literature there are different implementations of a Toffoli gate with the Clifford+T gate set, some optimizing its T-count, while there are others that optimize T-depth. The circuit in [68, 69] gives the lowest T-count of 4 and has a T-depth of 2. It uses logical AND gadget which does the multiplication. One advantage of this circuit is the fact that if we have a compute-uncompute pair then the uncomputation part does not require any T-gate, but it uses classical measurements. Another circuit is the one in [70] which has a T-count of 7, T-depth of 1 and uses 4 extra ancillae. Depending on whichever parameter we want to optimize, one implementation can be favoured over the other.
We have separated the cost after the computation of the encoding polynomials since the cost of this part can change depending upon the required operations, as discussed in Section 3.1.